OpenStudy (anonymous):

Need Help with 2 questions!

OpenStudy (babylove18):

what are they i can try and help?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Can not help sorry :'(

OpenStudy (amistre64):

what are your ideas?

OpenStudy (anonymous):

Would we combine like terms for the equation?

OpenStudy (anonymous):

Matrix :D

OpenStudy (amistre64):

no, just strip the equation of everything but the coefficients ...

OpenStudy (amistre64):

for the next one, add in the 'variables'

OpenStudy (amistre64):

3x-2y = 4 has the matrix [3 -2 4] why?

OpenStudy (anonymous):

you removed the coefficients?

OpenStudy (anonymous):

Variables*

OpenStudy (amistre64):

yep ... i stripped it down to the bare essentials

OpenStudy (anonymous):

so for number 1 would it be A?

OpenStudy (amistre64):

if we have ore than one equation to play with, we simply line up the like terms x -2y +5z = 1 4x + 3y = 3 -2x + z = 5 x y z c 1 -2 5 1 4 3 0 3 -2 0 1 5

OpenStudy (amistre64):

yes, A is the coefficient matrix for the system

OpenStudy (anonymous):

So i got it correct? Thanks.. lets move onto #2

OpenStudy (amistre64):

ive given a hint for #2, any ideas how we can apply it to 'undo' a matrix?

OpenStudy (anonymous):

no take me through lol no idea

OpenStudy (amistre64):

the options suggest we need variables of xyz, and some constant ... so, top it off with x y z c, and distribute them down the column

OpenStudy (anonymous):

would it be C?

OpenStudy (amistre64):

x y z c ----------- 1 -2 5 1 4 3 0 3 -2 0 1 5 1x -2y 5z =1 4x 3y 0z =3 -2x 0y 1z =5

OpenStudy (amistre64):

yes, C works out fine

OpenStudy (anonymous):

okay can you help me with a few more but they're not complicated ones like these, they're simple

OpenStudy (amistre64):

sure, why not

OpenStudy (amistre64):

simplest approach, trial and error, see what works and what doesnt

OpenStudy (anonymous):

im not sure would it be C or D for the first one, what do you think?

OpenStudy (amistre64):

neither, they dont fit the system ....

OpenStudy (anonymous):

aren't we combing like terms?

OpenStudy (anonymous):

how would i find the solution?

OpenStudy (amistre64):

not really, you are taking a point (a,b) and testing to see if it makes both statements true. i can tell right away that 11-8 is not 6 so D fails to be true for both parts

OpenStudy (amistre64):

8-4 is not 6, so C fails as well

OpenStudy (amistre64):

x+3y = 5 x+4y = 6 1 5 1+15 = 5 1+20 = 6 1,5 is not a solution either.

OpenStudy (amistre64):

if your objective is to determine the correct option, then work what is simplest for you. if your desire is to know how to work with a matrix ... then a longer approach is required.

OpenStudy (anonymous):

B?

OpenStudy (amistre64):

is 1,5 a solution to both equations?

OpenStudy (anonymous):

Im not typing anymore its just stuck on the screen lol.

OpenStudy (amistre64):

i think that is a client side affect, but yeah .... 'user' is typing seems to be a glitch

OpenStudy (anonymous):

Where are you getting 15 and 20 from?

OpenStudy (amistre64):

let x=1, and y=5 of course

OpenStudy (anonymous):

Lets start fresh move onto the 2nd one..

OpenStudy (amistre64):

simplest test is to let q=0 .. since that is an option

OpenStudy (amistre64):

do you know how to do matrix multiplication? inner product?

OpenStudy (anonymous):

yes matrix multiplication

OpenStudy (amistre64):

\[\begin{pmatrix}a&b\\c&d\end{pmatrix}\begin{pmatrix}x\\y\end{pmatrix}=\begin{pmatrix}n\\m\end{pmatrix}\color{red}{\implies}\begin{matrix}ax+by=n\\cx+dy=m\end{matrix}\] therefore the options given, some (x,y) will satisfy the system. test them out. or work it in whatever way is simplest for you. subsitution maybe? elimination? trial/error?

OpenStudy (amistre64):

i forgot to divide by a ...

OpenStudy (amistre64):

or by generalization of subsitution to get a formula if x=(n-by)/a c(n-by)/a + dy = m cn/a -bcy/a + dy = m y(d-bc/a) = m-cn/a y = (m-cn/a)/(d-bc/a) y = (am-cn)/(ad-bc)

OpenStudy (anonymous):

Would it be C? -1,0

OpenStudy (amistre64):

show me if C works or not

OpenStudy (amistre64):

im getting to much lag on the site .... C works for me. good luck